OpenStudy (anonymous):

Solve for k 1.6= 7.6 - 5(k + 1.1)

OpenStudy (anonymous):

I think the answer might be K= -2.3. I might be wrong.

OpenStudy (anonymous):

ok how did you get that

OpenStudy (anonymous):

alright so you have 1.6= 7.6-5(K + 1.1). First of all you have to get rid of parentheses. you need to do the distributive property. Since there is a minus sign beside 5, it would be negative five. You multiply -5 times K which gives you -5K. Then you have to multiply -5 times 1.1 which equals 5.5. That's step one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now your equation should be, 1.6= 7.6-5K + 5.5. Now you have to get rid of 5.5. To do that you would subtract 5.5. What you do to one side you must to to the other side of the equation. So on the left side of the equation, you would subtract 5.5, which gives you -3.9. That's step two.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Your equation should now look like this, -3.9=7.6-5K. Now we have to get rid of 7.6. To do that we must preform the inverse operation. We would subtract 7.6 on the right side, then on the left. This would give us -11.5= -5K. That's step three

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and then you would divide?

OpenStudy (anonymous):

Now we have -11.5=-5K. Now we have to get rid of the coefficient beside K. Now -5 and K are side by side, which represents multiplication. The inverse of multiplying is dividing. You would divide -5 by -5, and then do the same on the right side of the equation. This would give you -2.3=K, or K=-2.3. I'm not sure if this is right, so don't quote me on it. :)

OpenStudy (anonymous):

its right i checked it by substituting the variable with -2.3 thanks a lot

OpenStudy (anonymous):

You're welcome! :)